OpenStudy (anonymous):

Solve the system of equations by the substitution method. Choose one answer. a. (7, 2) b. (-7, 3) c. (6, 3) d. no solution

OpenStudy (turingtest):

post the system please

OpenStudy (anonymous):

Super sorry! =P

OpenStudy (anonymous):

first one : x = -3 + 5y plug it into the second : 9(-3+5y) -6y = 51 -27+45y-6y = 51 39y = 78 y=2 now use the first one : x = -3 + 5y to get : x = -3 +5*2 = -3 + 10 = 7 7,2 which means a